OpenStudy (anonymous):

You have to buy 26 tickets for a baseball game. You have $348 and must spend it all. How many of each type of ticket can you buy? (The front tickets are $18 and the back tickets are $12) pLease provide full explanation.

OpenStudy (anonymous):

$348 = front*$18 + back*$12 and front + back = 26

OpenStudy (anonymous):

with the money you have you could buy 29 x $12ticket and be within your savings, however, you are supposed to buy 26 tickets and spend it all so instead of having left over money, you should upgrade some of the tickets

OpenStudy (anonymous):

if you solve the system of equations, you get a collection of 26 tickets with the highest amount of front tickets that you can get :) $348 = front*$18 + back*$12 front + back = 26 use substitution or elimination to solve the system of equations which one do you prefer?

OpenStudy (anonymous):

Hey umm well i`m in grade six so can you please explain it in a somple easy way that I can understand. Thanks a lot.

OpenStudy (anonymous):

@mathessentials ^^

OpenStudy (anonymous):

your in 6th grade so you havnt learned system of equations yet. That is in Algebra 1, so don't worry about that yet

OpenStudy (anonymous):

Thanks but can you answer my question? You sound like a great person

OpenStudy (anonymous):

i can try. I am okay with systems of equations, so ill solve from where he left off. f=front b=back f18 + b12=348 f + b = 26 i am using elimination method btw. 18f+12b=348 -18f-18b=-468 -4b=-120 b=30 then plug 30 into one of the original equations f+b=26 f+30=26 f=-4 You cant have negative tickets so im not sure what to do there. Sorry, you should leave it blank and ask your teacher. Sorry i took so long i was hoping i would get the right answer but i guess i didnt. :p

OpenStudy (anonymous):

@sherry123 Verify for yourself that {back = 20, front = 6}